OpenStudy (anonymous):

How to SOLVE r^2-8r+16=0

OpenStudy (anonymous):

r^2 -8r + 16 = (r - 4)^2 which means you have to solve (r-4)^2 = 0 equals to (r-4) = 0 r = 4

OpenStudy (anonymous):

agree.......................:)